• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Areg

    Log entry time 09:08:58 on June29,2002

    Entry number 87048

    keyword=Shift summary -- Sat. owl

    2:00 Back to RCS
    2:00 cycled the rHRS quadrupoles, set the momentums to correct
    values. r-dipole is taking a while to stabilize.
    2:10 discovered that the control on the rHRS camera is broken.
    Called the tech.
    2:30 Access to the hall to reset the camera.
    2:55 Starting an optics run,1909.
    3:05 no beam.
    3:17 Starting collecting statistics for kin. RCS_A + radiator, run 1911.
    3:48 finishing run 1912
    3:58 Moved the spectrometers to RCS_B kin.
    4:06 Got current -- 50uA
    4:20 Finished run 1913.
    4:30 Roc14 crashed, rebooted.
    4:35 Beam down for 15 minutes.
    5:15 roc14 crashed
    5:55 Change to kinematics RCS_C.
    6:02 rHRS dipole finally stabilized...starting with run 1919.
    7:00 done with all the kinematics. requesting for access
    to change the scintillator for the s2 test.